solve this equation by substitution
7x-3y=-14
-3x+y=6

Answer by mando271(42)

Use either equation to solve for x or y. I'll use -3x+y=6 to solve for y cuz its the easiest.
-3x+y=6
y=3x+6
now i'll substitute this into the other equation to solve for x.
7x-3(3x+6)=-14
7x-9x-18=-14
-2x=4
x=-2
now plug this in to the solution for y
y=3(-2)+6
y=-6+6
y=0
(you can plug in both solutions into the original equations to see if it works)